OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 answer choices: 2 3 All real numbers No solution

OpenStudy (anonymous):

did you try and solve? if the left and right give the same expression, then it will be all real numbers would you like to try it ?

OpenStudy (anonymous):

I tried to solve it, and I got all real numbers

OpenStudy (anonymous):

well you can pretty much see that you are going to end up with \(-2x\) on both sides, so it will be all real numbers or no solutions

OpenStudy (anonymous):

Umm.. All real numbers?

OpenStudy (anonymous):

\[−2(x + 3) = −2(x + 1) − 4\] \[-2x-6=-2x-2-4\] \[-2x-6=-2x-6\] yup